satellite73 (satellite73):

\[\lim_{n\to \infty}\frac{3^n}{2^n+3^n}\]

OpenStudy (fibonaccichick666):

Do you know squeeze thm?

OpenStudy (anonymous):

i know it is one, i need a nice method

OpenStudy (aravindg):

How about taking 3^n common from denominatr

OpenStudy (aravindg):

nah ...wont work

OpenStudy (fibonaccichick666):

I'd just drop the 2^n off for sqeeze

OpenStudy (anonymous):

squueze is good since 3^n is always < 3^n + 2^n

OpenStudy (fibonaccichick666):

or divide everything by 3^n then simplify

OpenStudy (anonymous):

@FibonacciChick666 that looks good thanks

OpenStudy (christos):

go by derivatives d/(dn)2^-n

OpenStudy (fibonaccichick666):

np if it's a math theory course the dividing by 3^n is a better approach. for calc, I'd just use squeeze. :)

OpenStudy (anonymous):

not sure about squeeze though since \(\frac{3^n}{2^n+3^n} <\frac{3^n}{3^n}\) but what goes on the other sides?

OpenStudy (anonymous):

i would just drop off the \(2^n\) since it is not important, but still not sure how to use squeeze to get this

OpenStudy (fibonaccichick666):

for squeeze can it =0?

OpenStudy (anonymous):

no

OpenStudy (fibonaccichick666):

but that doesn't help hmmm

OpenStudy (anonymous):

if i you want to show it is one, you have to squeeze it between two things that are one

OpenStudy (fibonaccichick666):

could multiplying by the conjugate help?

OpenStudy (anonymous):

i like dividing by \(3^n\) top and bottom looks good to me

OpenStudy (fibonaccichick666):

haha yea it works, but now I want to know how to do it using squeeze lol

OpenStudy (fibonaccichick666):

btw how do you get your \[3^n\] in line with your text?

OpenStudy (anonymous):

you would have to find some other function \(g\) with \(g(n)<\frac{3^n}{2^n+3^n}\) and also \(\lim_{n\to \infty}g(n)=1\)

OpenStudy (anonymous):

i use \( instead of \[

OpenStudy (anonymous):

if you want to see any code, right click and choose "show math as" then "latex"

OpenStudy (fibonaccichick666):

thanks for that! and I'm thinking maybe \(e^{-x}\)

OpenStudy (fibonaccichick666):

but no that's zero too

OpenStudy (anonymous):

yeah you have to get something that goes to one not really sure what it would be

OpenStudy (fibonaccichick666):

hmmm... That is a head scratcher

OpenStudy (fibonaccichick666):

maybe \(\frac{x+1}{x-1}?\)

OpenStudy (fibonaccichick666):

yea, I think that works or it's inverse. I like the inverse better actually.

OpenStudy (e.mccormick):

Look at question #18 here: http://www.math.ucdavis.edu/~kouba/CalcOneDIRECTORY/liminfdirectory/LimitInfinity.html